# QQESPM: A Quantitative and Qualitative Spatial Pattern Matching Algorithm

### v1

@@ 6.2 Results
The average execution time by number of vertices in the search pattern was measured for each algorithm. The results, illustrated in Figure 4 (B), consistently demonstrate QQESPM’s superior runtime performance over the basic QQ-simple solution, regardless of the number of vertices. Notably, the observation that patterns with 5 or 6 vertices do not require longer execution times than patterns with 4 vertices can be explained by the increased complexity of search patterns, since in the same search area, patterns with too much keywords are less likely to have matches, and in these cases an early stopping of the query procedure can occur by identifying the non-existence of qq-n-matches in an early level.
The average memory allocation by QQESPM queries was also consistently lower for all dataset sizes and number of vertices evaluated, compared to the QQ-simple executions. The overall average memory allocation during queries was $284.16$ MB for QQESPM executions and $314.34$ MB for QQ-simple executions, highlighting the memory efficiency advantage of QQESPM over the QQ-simple trivial approach.

### v2

@@ 6.2 Results
The average execution time by number of vertices in the search pattern was measured for each algorithm. The results, illustrated in Figure 4 (B), consistently demonstrate QQESPM’s superior runtime performance over the basic QQ-simple solution, regardless of the number of vertices. Notably, the observation that patterns with 5 or 6 vertices do not require longer execution times than patterns with 4 vertices can be explained by the increased complexity of search patterns, since in the same search area, patterns with too much keywords are less likely to have matches, and in these cases an early stopping of the query procedure can occur by identifying the non-existence of qq-n-matches in an early level. The average memory allocation by QQESPM queries was also consistently lower for all dataset sizes and number of vertices evaluated, compared to the QQ-simple executions. The overall average memory allocation during queries was $284.16$ MB for QQESPM executions and $314.34$ MB for QQ-simple executions, highlighting the memory efficiency advantage of QQESPM over the QQ-simple trivial approach.